I don't understand how to solve or graph x-y=-2
y=-3

Answer by TimothyLamb(4379)

We have two linear equations:
---
L1: x - y = -2
L2: y = -3
---
To solve them means to find the point or points of intersection, if any, between L1 and L2.
---
Substitute -3 (the value of y from L2) for y in L1:
---
L1: x - y = -2
L1: x - (-3) = -2
---
Now solve for x:
---
x - (-3) = -2
x + 3 = -2
x = -2 - 3
x = -5
---
Now plug-in (-5) for x into L1 and solve for y:
---
L1: x - y = -2
L1: (-5) - y = -2
(-5) - y = -2
y = (-5) + 2
y = -3
---
Answer:
L1 and L2 intersect at the point p0 = (-5, -3)
---
To check the answer, plug-in the point p0 into L1 and into L2.
The point p0 satisfies both L1 and L2 so its a correct answer.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php